OpenStudy (anonymous):

Please Help Ill Give Medal and Fan U!!!!!!! Stacey is studying the sale of a particular brand of cereals from the year 1995 to 2004. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t)= t^2 +6t+48 Part A: What does the y-intercept of the graph of the function represent? Part B: What is the reasonable domain of the graph of the function? Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year?

OpenStudy (anonymous):

The y-intercept will be the +48 at the end of the function, so that takes care of Part A.

OpenStudy (anonymous):

What do u mean by +48

OpenStudy (anonymous):

The function is S(t)= t^2 +6t+48, the y-intercept will always be the last number if it's in that form.

OpenStudy (anonymous):

+48 is at the end the '+' just indicates that the '48' is a positive number, and not a negative.

OpenStudy (anonymous):

ohhh ok thanks.

OpenStudy (anonymous):

(That's Part B)

OpenStudy (anonymous):

Sorry let me correct that..

OpenStudy (anonymous):

that's fine go ahead.

OpenStudy (anonymous):

The domain should be 0<t<183, because Stacey is studying between the years 1995-2004, which is 9 years. On your graph S(t)= t^2 +6t+48, when t(years) = 9, y = 183. So 183 would be your highest, and 0 would be your lowest, because you cannot have negative sales of cereal.

OpenStudy (anonymous):

We're dealing with 't' not 'x' for the variable.

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

To find the average rate of change from year 1 to year 4, just plug in 1 and 4 for t, in the equation: S(t) = t^2 + 6t + 48 S(1) = (1)^2 + 6(1) + 48 S(1) = 1 + 6 + 48 S(1) = 7 + 48 S(1) = 55 So year 1 = 55. S(t) = t^2 + 6t + 48 S(4) = (4)^2 + 6(4) + 48 S(4) = 16 + 24 + 48 S(4) = 40 + 48 S(4) = 88 So year 4 = 88 Now we can plug in our values (1, 55) and (4, 88) into the slope equation: m = y2 - y1 / x2 - x1 y2 = 88 y1 = 55 x2 = 4 x1 = 1 m = 88 - 55 / 4 - 1 m = 33 / 3 m = 11 So the average rate of change from year 1 to year 4 is 11.

OpenStudy (anonymous):

And that's Part C.